OpenStudy (anonymous):

an=n^3/(n-1)!....write the first four terms of the sequence whose general term is given?...please help

OpenStudy (midhun.madhu1987):

put n = 0, 1, 2,3 and 4 in the general equation...

OpenStudy (cwrw238):

n = 1,2,3,4 for first term we get 0! as the denominator do you know value of 0! ?

OpenStudy (anonymous):

is it 3?

OpenStudy (cwrw238):

0! = 1 so first term = 1^3 / 1 = 1

OpenStudy (cwrw238):

2nd term = 2^3 / (2-1)!

OpenStudy (cwrw238):

then plug in 3 and 4 into the general formula

OpenStudy (anonymous):

Am getting 1/12, but am not sure am right

OpenStudy (anonymous):

sorry, that was a different question

OpenStudy (anonymous):

1/0, 8/0, 27/2, 32/3 ..... is that correct?

OpenStudy (cwrw238):

no as i said 0! = 1 so first term is 1^3 / 1 = 1 second = 2^3 / 1! = 8/1 = 8

OpenStudy (cwrw238):

do you now what n! means? pronounced n factorial

OpenStudy (anonymous):

number? or general term

OpenStudy (cwrw238):

its best explained by an example 3! = 3*2*1 6! = 6*5*4*3*2*1 - just keep multiply down to the number 1

OpenStudy (cwrw238):

so the 3rd term is 3^3 / (3-1)! = 27/2! = 27 / 2*1 = 27/2 or 13.5

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

ok...am now getting it

OpenStudy (anonymous):

Thanks a million

OpenStudy (cwrw238):

so can you work out the 4th term?

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

32/3 right?

OpenStudy (cwrw238):

4^3 / (4 -1)! = 64 / 3! = 64 / 6 = 32/3 yes

OpenStudy (cwrw238):

gotta go right now

OpenStudy (anonymous):

wow... now i get it

OpenStudy (anonymous):

byee, thanks for your help